OpenStudy (anonymous):

Please someone!!! Find a linear quadratic and a cubic model for the data. Which models best fits the data. X=3 8 15 21 Y=7 11 26 44

jimthompson5910 (jim_thompson5910):

This is a real pain to do by hand, so the next best thing is either a calculator or something like excel

OpenStudy (anonymous):

i need to show some work though

jimthompson5910 (jim_thompson5910):

are there any examples in your book that show how to show the work hopefully there's at least a formula shown or something because I'm blanking on the formula needed...one sec

jimthompson5910 (jim_thompson5910):

if you wanted to do linear regression, then you would use the formulas on this page http://serc.carleton.edu/introgeo/teachingwdata/StatRegression.html

OpenStudy (anonymous):

you know what I read my book and it says to use a calculator so i hope my teacher is ok with little work. can u help me with this one though! Find a quadratic function whose only real zeros are 4 and 6

jimthompson5910 (jim_thompson5910):

the formulas for quadratic regression can be found here http://www.azdhs.gov/lab/license/tech/calibration-training/calibration/documents/12_Quadratic-Least-Squares-Regression-Calib.pdf ** look at page 10 ** but as you can see it gets really ugly (if it wasn't ugly before) really fast...

jimthompson5910 (jim_thompson5910):

Find a quadratic function whose only real zeros are 4 and 6 x = 4 or x = 6 x - 4 = 0 or x - 6 = 0 (x - 4)(x - 6) = 0 x^2 - 6x - 4x + 24 = 0 x^2 - 10x + 24 = 0 So the quadratic with the two real zeros 4 and 6 is y = x^2 - 10x + 24

OpenStudy (anonymous):

thanks so much u are a life saver!!!

OpenStudy (anonymous):

The cubic looks real good. The Mathematica calculation is attached.